OpenStudy (anonymous):

URGENT!! factor the following expression 20xy

OpenStudy (anonymous):

in terms of what?

OpenStudy (anonymous):

what do you mean

OpenStudy (icalibear):

10 * 2 = 20 5 * 2 = 10

OpenStudy (anonymous):

what are you factoring out?

OpenStudy (zehanz):

In 20xy = 20*x*y, there are already 3 factors: 20, x and y. x and y are unknown, so they could be any number. It is therefore not possible to factor them any further. The factor 20, on the other hand, is composite, i.e. it is not a prime number, but it is the (unique) product of prime numbers. As @icalibear already pointed out, 20 =2*2*5, so imho the answer to your question is: \[2^2*5*x*y\]